### v1

# Comprehensive benchmarking of large language models for RNA secondary structure prediction

Abstract
Inspired by the success of large language models (LLM) for DNA and proteins, several LLM for RNA have been developed recently. RNA-LLM uses large datasets of RNA sequences to learn, in a self-supervised way, how to represent each RNA base with a semantically rich numerical vector. This is done under the hypothesis that obtaining high-quality RNA representations can enhance data-costly downstream tasks. Among them, predicting the secondary structure is a fundamental task for uncovering RNA functional mechanisms. In this work we present a comprehensive experimental analysis of several pre-trained RNA-LLM, comparing them for the RNA secondary structure prediction task in an unified deep learning framework. The RNA-LLM were assessed with increasing generalization difficulty on benchmark datasets. Results showed that two LLM clearly outperform the other models, and revealed significant challenges for generalization in low-homology scenarios.
Introduction
Ribonucleic acid (RNA) plays a crucial role in many fundamental biological processes, such as gene expression, cell signaling, and post-transcriptional regulation [20, 56]. As in proteins, RNAs function and interaction with other molecules are deeply related to their structure. For example, determining RNA structure is essential for RNA-based therapeutics such as mRNA vaccines [34]. Among all RNA transcripts, only 5% is responsible for protein coding. At the same time, a very large remaining portion is non-coding RNA (ncRNA) [7], which in many cases adopt specific structures to perform important biological functions [6]. Experimental results show that, to some degree, sequence determines the secondary structure, and thus, their function [33, 16, 52]. Therefore, the long-established and yet unresolved RNA secondary structure prediction is major challenge in computational biology today [4, 5].
During many years, experimental molecular structure technologies like nuclear magnetic resonance, X-ray crystallography and cryogenic electron microscopy have produced several RNA structures [19, 45]. However, despite the large number of ncRNA sequences available, most of their structures and functions remain still unknown [57]. At the same time that large amounts of unlabeled RNA sequence data were produced by high-throughput sequencing technologies, pre-trained RNA language models have started to be used for modeling the semantic space of RNA sequences, intending to facilitate its understanding. Motivated by the success of large language models (LLM) in proteins [32, 37, 18, 3, 2, 39, 14, 31] and DNA [26], several RNA LLM have recently appeared [1, 36, 8, 55, 61, 58, 54] with the potential to be used for improving several important RNA-related tasks, among which one of the most relevant is RNA secondary structure prediction.
RNA-LLM attempt to effectively embed RNA bases using deep representation learning, in particular the one developed in the field of natural language processing [18, 39]. Most RNA-LLM are based on bidirectional encoder representations from transformers (BERT), which was designed to generate context-sensitive distributed test token (word) representations [12]. The relevance of LLM for transfer learning from sequences to downstream tasks have been analyzed in detail for proteins [25, 53]. In the case of RNA, the hypothesis is that LLM representation of nucleotide composition and sequence motifs can help characterize the structure and function of a sequence, analogously to how the meaning of a sentence is determined by the grammatical structure of natural language. Thus, word embedding techniques for natural language have been applied to bases for RNA sequences, obtaining a wide variety of BERT-based architectures trained on different databases. However, in spite that some theoretical reviews on the applications and utility of LLM in bioinformatics have recently appeared [26, 50, 35], to the best of our knowledge all the available RNA-LLM have not yet been comparatively evaluated for the secondary structure prediction task, considering a fair experimental setup, with the same datasets and structure prediction model. Since LLM are pre-trained models, it is important to analyze how they were trained to assess generalization capabilities on new sequences/structures, considering homology-aware partitions and cross-family predictions.
Table 1: RNA large language models included in this study.
RNA-LLM Year Emb. Dim. Pre-train seqs. Databases Architecture (layers) Parameters Repository
RNABERT[1] 2022 120 76,237 RNAcentral Transformer (6) 509,896 [Uncaptioned image]
Pre-training with masked language modeling (MLM)[13] masking 15% random bases and predicting from neighboring bases, with data augmentation taking 10 copies of each ncRNA and applying 10 different mask patterns. Structural alignment learning (SAL) pre-training task based on seed alignment for each RNA family with Rfam. Token embedding randomly generated 120 dimensions RNA bases. Positional embedding for each base in the sequence. Stack of transformers with multi-head self-attention and feed-forward neural network. Softmax output layer. Cross-entropy loss function.
RNA-FM [8] 2022 640 23,700,000 RNAcentral Transformer (12) $\sim$100,000,000 [Uncaptioned image]
Pre-training with MLM and 15% random masking. Based on BERT original architecture. Transformer encoder blocks, with multi-head self-attention modules and feed-forward layers. Softmax output layer. Cross-entropy loss function.
RNA-MSM [61] 2024 768 3,087,138 Rfam Transformer (12) $\sim$96,000,000 [Uncaptioned image]
Multiple sequence alignment (MSA) inspired by AlphaFold2 [21]. Homology search and MSM of training sequences with RNAcmap3 database [9] as augmented input. Two modules: embedding and MSA transformer. Initial embedding layer and two learnable position-embedding layers. Stack of MSA transformer blocks with a residue and sequence attention layer followed by a feed-forward layer with GELU activation functions. Cross-entropy loss function.
ERNIE-RNA [58] 2024 768 20,400,000 RNAcentral Transformer (12) 86,000,000 [Uncaptioned image]
Pre-training with MLM and 15% random masking. Based on the Enhanced Representation through Knowledge Integration (ERNIE) framework [47]. Modified BERT that incorporates base-pairing informed attention bias when calculating attention. From the second layer onward, the bias of each layer is determined by the attention map of the previous layer. Transformer blocks with attention heads followed by a feed-forward layer. Cross-entropy loss function.
RNAErnie [54] 2024 768 $\sim$23,000,000 RNAcentral Transformer (12) $\sim$105,000,000 [Uncaptioned image]
Pre-training with MLM and 15% random masking. Motif-aware additional pre-training strategy involving motif-level random masking from motif databases [24], and random subsequence masking. Built upon the ERNIE framework. Transformer blocks followed by a feed-forward layer. Cross-entropy loss function.
RiNALMo [36] 2024 1280 $\sim$36,000,000 RNAcentral + Rfam + Ensembl Transformer (33) $\sim$650,000,000 [Uncaptioned image]
Pre-training with MLM and 15% random masking and FlashAttention-2 [11]. BERT-style encoder-only transformer. Position encoded using rotary positional embedding [46]. Each block comprises a multi-head attention of 20 heads. Feed-forward network with two linear layers and SwiGLU activation functions [41]. Residual connections with layer normalization to stabilize training. Cross-entropy loss function.
In this work, we provide a comprehensive experimental comparison of the latest pretrained RNA-LLM, using a unified and consistent experimental setup. It offers an independent, third-party assessment following the best practice guidelines for bias-free evaluations in the community [60]. We provide a description of each model, together with an experimental testing in four benchmarks of increasing complexity, with the corresponding source code and datasets to ensure reproducibility, supporting the development and boosting of future improvements on RNA secondary structure prediction based on LLM.
Results
Self-supervised RNA-LLM and prediction model
In the last three years, a number of RNA-LLM have appeared in literature. We summarize in Table 1 the most recent RNA-LLM, included in this study according to their availability as open source tools. All models and pre-training methods are detailed in the Methods section.
For the comparative analysis, embeddings of each RNA-LLM feed the same deep learning architecture (Figure 1a-e) for RNA secondary structure prediction. The rationale behind this is that it should not be necessary to optimize the classifier hyperparameters for each LLM, since they have complex architectures and have already been trained with a large amount of data. Thus, according to the transfer learning paradigm, the embeddings will have enough information to solve the downstream task. Moreover, any difference measured in performance will be due to the LLM and not because of the classifier architecture. This is the only trainable part of the experimental setup, which uses the train-test partitions defined for each benchmark dataset (detailed in Methods).
Figure 1: |||| RNA-LLM embeddings and deep neural architecture of the prediction model. a, Flow diagram for RNA secondary structure predictions. Each LLM was downloaded from its official repository, frozen and used to get per-nucleotide embeddings for each sequence in the benchmark datasets. The embeddings go through a fully connected layer, outer concatenation, two 2D ResNet blocks and a final 2D convolution. This flow is explained in detail in aligned panels b-e. b, One-hot encoding, of size $4\times L$, and LLM representation with size $d\times L$, where $d$ is the embedding dimension and $L$ is the padded sequence length. A fully connected layer reduces the input dimension to $M/2$. c, The $M/2\times L$ projection is transformed to a $M\times L\times L$ tensor using outer concatenation. The position $(i,j)$ contains the concatenated representation of nucleotides $i$ and $j$. d, Then, there are two 2D ResNet blocks with 2D Convolution, Instance Normalization and ReLU. ResNet Blocks 1 has kernel size 1 and ResNet Blocks 2 has kernel size 3. e, A final convolution yields the output scores as a $L\times L$ connection matrix. f-l, RNA-LLM embeddings projected with UMAP for dimensionality reduction. The RNA families of the ArchiveII dataset are highlighted with different colors. Sequences within the same RNA family are expected to be close in the dimensionally reduced space.
LLM separate structural families even when they were not pre-trained with family information
A useful RNA embedding for the RNA secondary structure prediction task is expected to well-represent not only sequence properties, but also structural aspects. Therefore, RNA embeddings encoding similar features should share a common region in the multidimensional embedding space. To have a preliminary insight into this topic, a qualitative comparison of LLM embeddings of the ArchiveII dataset was performed by nonlinearly reducing them to three dimensions using the uniform manifold approximation and projection (UMAP) [30]. After that, each RNA sequence was depicted with a different color according to its corresponding RNA family (Fig. 1f-l).
The one-hot encoding (Fig. 1f) was used as reference. In this case, as for RNABERT (Fig. 1g) representations, there is a large overlap and mixing of sequences that belong to different RNA families; that is, in the same close region there are sequences that are very different in length, structure, and function. Instead, in the case of RNA-FM (Fig. 1h), ERNIE-RNA (Fig. 1j) and RiNALMo (Fig. 1l) each RNA family is mostly separated from the others. The 5s family (gray), which is the largest family, is distributed in several regions of the projected space. In the case of RNA-FM, the 5s, tmRNA (red) and tRNA (pink) families are well-separated among them. Overall, it can be stated that for RiNALMo and ERNIE-RNA, most RNA sequences and families are well-separated in the 3D projected space. In the case of RNA-MSM (Fig. 1i), there are two large and separated clusters but with a mix of families within. One large cluster has tmRNA, RNaseP (yellow), srp (brown), 23s (green), and telomerase (violet); while the other cluster has 5s, srp and tRNA. RNAErnie has a mix of very well-separated and cohesive groups of sequences belonging to the families tRNA, srp and 5s; but there is also a large overlap among sequences of the other families. The telomerase family, the longest one, can hardly be distinguished in any case, being cohesively and separately represented only in the case of RiNALMo and ERNIE-RNA.
Performance on homology-challenging datasets
Figure 2 shows the comparative results among the RNA-LLM here reviewed in terms of $F_{1}$ violin plots. In each plot the results are presented in order, from best (left) to worst (right) median $F_{1}$. Figure 2a shows the results for 5-fold random partitions on the ArchiveII dataset. It can be seen that most predictors achieve a median performance above $F_{1}=0.60$, except only for the one based on one-hot encoding ($F_{1}=0.57$). The top-3 RNA-LLM achieve a very high performance ($F_{1}>0.90$). Those are ERNIE-RNA and RiNALMo, both with $F_{1}=0.95$, and RNA-FM with $F_{1}=0.91$. RNAErnie and RNA-MSM achieve an intermediate performance, $F_{1}=0.76$ and $F_{1}=0.74$, respectively. Finally, RNABERT is the RNA-LLM with the lowest average performance ($F_{1}=0.62$). For this random $k$-folding, all RNA-LLM except RNABERT achieve higher performance than the classical prediction method (LinearPartition-C [59]).This result shows that the neural architecture used in the classifier has the capability of obtaining high performance in predictions. However, it is well-known that the sequence homology between train and test at random partitions favors methods based on deep learning; thus performance tends to be overly optimistic.
Figure 2: |||| Comparative results among RNA-LLM on the RNA secondary structure prediction task for different benchmark datasets of increasing complexity. Each method has a different color and the performance of LinearPartition-C is shown for reference with a horizontal blue line in all datasets. Next to RNA-LLM performances, sequence length distribution for each dataset is shown in blue. a, ArchiveII 5-fold random cross-validation. A Wilcoxon test for paired samples with Bonferroni correction indicates that all differences are statistically significant ($P<0.0001$, $N=3,864$). b, bpRNA train-test partitions with controlled homology. All differences are statistically significant ($P<0.0001$, $N=1,305$) except for one-hot and RNABERT. c, bpRNA-new dataset, for RNA families not seen during training. All differences are statistically significant ($P<0.0001$, except for RNAErnie versus RNA-MSM and one-hot with $P<0.05$, $N=5,401$). d, PDB-RNA dataset, with RNA sequences extracted from PDB. Most differences are not statistically significant among RNA-LLM, except for RiNALMo and ERNIE-RNA with respect to RNA-FM, RNABERT and RNAErnie. Details of statistical analysis in Supplementary Fig. S1a-d. e, Average $F_{1}$ per motif type on bpRNA-new. f, Performance accounting only non-canonical interactions on PDB-RNA.
For a deeper analysis, we have extended the experiments on standard datasets having controlled levels of homology between training and testing partitions. Figure 2b shows the test $F_{1}$ for bpRNA dataset, with a partition at 80% sequence-identity cutoff between training (TR0+VL0) and testing (TS0) (see Methods). The results show that now, in a harder testing setup, most methods lowered performance in comparison to random partitions. Here, again, ERNIE-RNA and RiNALMo have both the best median scores in the test set, but in this case with a large drop in performance: $F_{1}=0.68$ and $F_{1}=0.64$, respectively. These two RNA-LLM are the only ones that achieve better performance than the classical prediction method ($F_{1}=0.60$). With median performance below it, RNA-FM has $F_{1}=0.52$, RNA-MSM has $F_{1}=0.42$, and RNAErnie has $F_{1}=0.41$. Both one-hot and RNABERT reached similar and very poor results, $F_{1}=0.35$ and $F_{1}=0.33$, respectively.
Figure 2c shows the results of testing already trained models with the bpRNA-new dataset, which was designed to evaluate the prediction on new RNA families, never seen during training. The performance of all methods has been noticeably reduced, being the median of the RNA-LLM in all cases below the classical folding method ($F_{1}=0.71$). The only model with performance very close to the classical one is ERNIE-RNA, having a median $F_{1}=0.67$ (with large dispersion). In all the other cases, the median $F_{1}$ is below $0.45$. Additionally, note that the one-hot encoding moved up one more position in the ranking, slightly outperforming RNAErnie in this test set. This large dataset was used to analyze the prediction scores per structural motif, as seen in Fig. 2e. It can be seen that most models behave similarly for each motif type, with ERNIE-RNA showing a higher performance in stem base-pair predictions. Overall, ERNIE-RNA, RiNALMo and RNA-MSM achieved the best scores for each motif.
Results on a more challenging dataset are shown in Fig. 2d, where every RNA-LLM was tested in a set of 62 RNA sequences whose secondary structures were inferred from high-resolution RNA X-ray 3D structures from the PDB dataset. Results clearly show a hard fall in results for all the RNA-LLM, with median performances below $F_{1}=0.40$, lower than the classical folding method, which achieved a moderate performance of $F_{1}=0.60$. Notably in this dataset, the one-hot encoding is within the top-3 methods, outperforming RNA-FM, RNAErnie and RNABERT; with almost the same performance that RNA-MSM, ERNIE-RNA and RiNALMo. It can be said that, for this prediction challenge, no RNA-LLM provides significant improvements of transfer learning over a one-hot encoding. This dataset also contains non-canonical interactions, although most of the methods fail to reach a median $F_{1}$ score higher than $0.2$ (Fig. 2f), with a slightly better performance for ERNIE-RNA and RiNALMo.
Cross-family benchmarks
For assessing inter-family performance, a family-fold cross validation in the ArchiveII dataset was performed. That is, one family was left out for testing per cross-validation fold, and the rest of the families were used for training. This eliminates most of the homology to the training set, providing a hard measure of performance and, thus, allowing estimating future performance on novel RNAs that do not belong to any known family. These results are shown in Fig. 3, along with the distributions of sequence lengths and minimum structural distances [38] between each test sequence and all the sequences in its corresponding training fold.
Figure 3: |||| Inter-family structure prediction based on RNA LLM. Performance evaluated on the 9 RNA families of the ArchiveII dataset. Each boxplot represents the $F_{1}$ performance of all methods for a given family in the test set. The median performance of LinearPartition-C is shown, for each family, as a horizontal blue line. Below RNA-LLM performances, sequence length distribution for each dataset is shown in blue and distribution of minimum test-train structural distance [38] is shown in orange. a, tRNA family. A Wilcoxon test for paired samples with Bonferroni correction indicates that all differences are statistically significant ($P<0.0001$, $N=557$), except in the case of RiNALMo with ERNIE-RNA, and RNABERT with one-hot. b, 5s family. All differences are statistically significant ($P<0.0001$, $N=1283$). c, tmRNA family. All differences are statistically significant ($P<0.0001$, $N=462$). d, RNaseP family. Only ERNIE-RNA with RiNALMo, RNA-MSM with RNAErnie, and one-hot with RNABERT are not statistically significant. e, srp family. Differences for the top 4 methods are statistically significant ($P<0.0001$, $N=918$). f, grp1 family. Differences for the top 2 methods are statistically significant ($P<0.0001$, $N=74$). g, 23s family. Top 2 methods are significantly better than the rest of all the methods ($P<0.01$, $N=15$). h, 16s family. RiNALMo and ERNIE-RNA are statistically different and are also statistically different from the rest of the methods ($P<0.0001$, $N=66$). i, telomerase family. ERNIE-RNA and one-hot are statistically different from RNABERT, RiNALMo, RNA-MSM and RNAErnie. Details of statistical analysis in Supplementary Fig. S2.
In the case of the tRNA family (Fig. 3a), the one with the shortest sequence length as shown in the panel below, most methods achieve performance above $F_{1}=0.40$. RNABERT and one-hot achieved both the same performance here, close to $F_{1}=0.46$. The four best methods for this family are RiNALMo ($F_{1}=0.86$), ERNIE-RNA ($F_{1}=0.84$), RNAErnie ($F_{1}=0.81$) and RNA-FM ($F_{1}=0.79$). In this case, which can be considered the easiest one for a cross-family validation due to the small length of sequences and test/train structural distance below 0.5, four of the RNA-LLM outperform the classical LinearPartition-C, which achieved $F_{1}=0.75$. For the 5s family (Fig. 3b), the results for all RNA-LLM are very similar to the previous family, again due to the length of the sequences (the second mean shortest family) and low structural distance between this family and all the remaining families in the training set. In this case, one-hot representation, RNABERT and RNA-MSM are below $F_{1}=0.26$, while RNAErnie and RNA-FM achieve performances around $F_{1}=0.55$, and again ERNIE-RNA and RiNALMo achieved both the best results, with $F_{1}=0.84$ and $F_{1}=0.82$, respectively, above the classical method ($F_{1}=0.78$). For the tmRNA (Fig. 3c) and RNaseP (Fig. 3d) families, the conclusions are very similar as before. Although ERNIE-RNA and RiNALMo have performance 10% lower than with the previous families, in these cases achieved even higher performance than the classical method. In the case of the tmRNA family (Fig. 3c), the classical method has $F_{1}=0.38$, while ERNIE-RNA and RiNALMo have almost twice the performance with $F_{1}=0.73$. In the case of the RNaseP (Fig. 3d) family, those two RNA-LLM methods are more than 10% better than the classical one. Here, it is important to note that these two families are very similar to each other in average sequence length, around 300 nt, but 3 times longer than tRNA and 5s families.
In the case of the srp family (Fig. 3e), there is an important change in the length distribution, which becomes bimodal, and in the structural distances distribution, which clearly exceeds the 0.5 mark. Here, the classical method, ERNIE-RNA and RiNALMo achieve extremely close results while all other RNA-LLM are far below them. ERNIE-RNA has median $F_{1}=0.67$ while the classical method achieved $F_{1}=0.68$. Notably, in the case of the grp1 family (Fig. 3f), which has larger average length size than srp, RiNALMo is the only LLM that slightly outperforms the classical method ($F_{1}=0.64$ versus $F_{1}=0.61$). In the case of the 23s and 16s families (Figs. 3g and 3h, respectively), it can be seen that the one-hot representation, RNABERT, RNA-FM, RNA-MSM and RNAErnie embeddings are all below $F_{1}=0.25$, a very low performance. However, ERNIE-RNA and RiNALMo achieve almost twice that performance with median values near to $F_{1}=0.60$, and the classical method achieves slightly better results ($F_{1}=0.69$ for 23s and $F_{1}=0.66$ for 16s). Notably, this last family has a wide variety of sequence lengths and structural distances. Finally when the telomerase family (Fig. 3i) is used in the test set, all RNA-LLM achieve extremely poor results below $F_{1}=0.15$. This family has the longest sequences. Here the classical method has a large advantage over all RNA-LLM, although it is low compared to the other families.
Discussion
We benchmarked 6 RNA-LLM for their ability to predict RNA secondary structure structure based only on the sequence, in 4 datasets of increasing complexity. First we performed a visual comparison of the LLM embeddings with a UMAP projection, analyzing the distribution and separation of the RNA families. We found that RiNALMo and ERNIE-RNA were the models that could better represent and separate the RNA families in the projection without almost overlap. Notably, as Table 1 shows, those RNA-LLM are the biggest ones, pre-trained with the largest and most varied datasets.
We tested each RNA-LLM with exactly the same experimental setup and deep neural network architecture for RNA secondary structure prediction in several benchmark datasets, increasing the difficulty from simple random partitions to very low homology between training and testing partitions. Overall, RiNALMo and ERNIE-RNA maintained an acceptable level of generalization capability even in the hardest test, providing equivalent performance to the classical folding method. Remarkably, in that hardest scenario the trivial one-hot representation achieved a performance similar to the best performing RNA-LLM. Thanks to the transfer learning paradigm, one would have expected that LLM would help to generalize even with a small training set and prediction model, because all the relevant information was learnt in the embeddings. That was not the case and it was probably an indication that there is still important information that cannot be fully captured even by the LLM pre-training. The results achieved might be explained by the fact that the structures of sequences in the testing sets were really very different from those seen during pre-training, or they were a small minority, thus the generalization capability of the prediction model could not be really benefited from using a LLM representation.
For the most difficult secondary structure prediction task, we removed from training all the sequences of one specific testing family. In this setup, most of the RNA-LLM achieved a high performance for the two shortest families. However, as the average length of the sequence to be predicted increased, the performance of most RNA-LLM lowered below $F_{1}=0.50$. Overall, in 4 out of 9 cases RiNALMo and ERNIE-RNA outperformed the classical method. In two cases no differences were observed, and in the last three families the classical method was the best. In the case of the telomerase RNA family, the most different from the other ones, with very few samples in all datasets and the largest average sequence length, the results were all poor, up to three times below the classical method. In any case, in this RNA family the classical method also obtained a performance without practical usefulness, barely reaching $F_{1}=0.50$.
Our study showed that RNA-LLM in combination with a deep neural network for prediction were surprisingly capable of outperforming a classical method in half of the RNA families, even though cross-family prediction was a task historically dominated by thermodynamic methods. Although in this study we have not delved into the analysis of classical methods, because those are outside the scope of the RNA-LLM comparison, it is important to note that their historical superiority is due to a limited understanding of the concept of data-driven learning. This is intuitively associated only to machine learning methods, but all thermodynamics-based methods also have data-driven learning. For machine learning-based methods, humans build a software with rules to be trained automatically from domain examples, thus excluding some of those examples for testing is very simple. Conversely, for classical methods humans learn themselves from data samples, and then build a software with those learned rules. That is: they observe many sequences and structures for years, test and improve their models based on those observations, and even incorporate measurements of real structures as parameters of their models [29]. In these cases, it is much more difficult to simply exclude training samples from testing sets. This is why a new methodology is required to make a fairer comparison with thermodynamic methods.
We found that those RNA-LLM trained in the self-supervised stage with the largest and most varied number of sequences, and with the largest number of trainable parameters (ERNIE-RNA and RiNALMo), were also those that better represented the different RNA families in the projected UMAP space and those that accordingly achieved higher performance overall, very consistently and in all cases. At the same time, our experimental design clearly revealed the most important challenges that remain unsolved in RNA secondary structure prediction yet. We generated the first systematic benchmarking for this prediction task with LLM. Our experimental methodology, the curated datasets and the full source code are a key tool for analysts to navigate the space of available RNA-LLM methods, and constitutes a reference base for developers towards building more efficient prediction methods. We are confident that the criteria and analysis processes defined here can become a benchmark for future systematic investigations of RNA-LLM performance.
Acknowledgments
This work was supported by Argentine National Scientific and Technical Research Council (CONICET), ANPCyT (PICT 2022–0086) and UNL (CAI+D 2020 115). We also acknowledge the support of NVIDIA Corporation for the donation of GPU used for this research.
Author contributions
GS and DHM conceived the idea, designed and initiated the study. DHM coordinated collaborations and communication between team members. LZ, LB and DHM developed the experimental design and source code. GS wrote the first manuscript. GS and DHM edited the manuscript with contributions from all coauthors. DHM and MG created figures. LZ, LB and DHM worked on data visualization. LZ and LB implemented the deep learning model and built the repositories. LB and LZ curated and managed the datasets. LDP, GS and DHM performed statistical analyses. All the authors perform computational evaluation of models. All authors reviewed and approved the final manuscript.